OpenStudy (anonymous):

Salma drove 250 miles using 9 gallons of gas. At this rate, how many gallons of gas would she need to drive 275 miles

OpenStudy (anonymous):

you can get the average gallons per mile with the existing data: DROVE: 250 mi TANK DECREASE: 9 gallons

OpenStudy (anonymous):

So.. would she need 10? Or 9 point something..

OpenStudy (anonymous):

the total distance (100%) is 275 mi we have given gallon usage for 250 of 275 miles \[9 gallon = 250 mi\]

OpenStudy (anonymous):

you can divide by 250 mi on both sides: \[9/250mi gallon = 1 \] and then multiply by 275 mi \[9 \times 275mi/250mi = 275 mi\]

OpenStudy (anonymous):

Travel is 275 / 250 times longer, usage is also 275/250 times higher.

OpenStudy (anonymous):

I got 9.9.. Is that correct?

OpenStudy (mathstudent55):

Use a proportion: 250 miles is to 9 gallons as 275 miles is to x gallons "is to" means a ratio. 250:9 = 275:x A ratio can also be written as a fraction: \( \dfrac{250}{9} = \dfrac{275}{x} \) Now cross multiply and solve for x. 250x = 9 * 275 250x = 2475 x = 9.9 Answer: 9.9 gallons. You are correct!

OpenStudy (anonymous):

I get the same result :) we can check that the result makes sense: 275 = 250 + 25 so we get an additional 10% in distance (25 of 250 = 10%) and the gallons are also increased 10% (which is +.9 of 9 gallons)